# Towards Generalised Pre-Training of Graph Models

### v3

@@ 3.1 Hypothesis 1: Excluding Features
Supporting Hypothesis 1 requires producing a model that has learnt useful representations of graphs in general. Our first step to achieving this is to construct a dataset that covers a variety of graph domains. Such a dataset should contain graphs that present a large diversity of multi-node patterns. Our approach is simple: include graphs from several domains and assume that there is adequate coverage of the multi-node patterns that occur in real-world graphs for a given set of downstream tasks. The same approach has been taken in the domains of text [3] and images [24], where enormous numbers of data samples are scraped from the web for generative pre-training. This relies on two assumptions 1) this dataset adequately covers the relevant space for downstream transfer, and 2) a single model can usefully represent such a space. These assumptions, as in other fields, are justified through performance metrics on downstream tasks on seen and unseen domains.
@@ 3.2 Hypothesis 2: Sample Diversity
From Hypothesis 2 we expect that for a given task models trained on a more diverse pre-training domain can bring better transfer than a model trained on the target domain. We do not argue that this is always true, only that in some settings out-of-domain samples can be more useful in pre-training than in-domain. This has clear parallels to the performance-by-scale approach taken in other fields [28, 6, 23]. To this end we compare two models, the first trained on a dataset from a single domain, as is the current practice. This single domain matches that of the downstream target dataset(s). The second model is instead trained on the rest of the data while excluding the target domain.
@@ 4 Experiments
Our experiments are guided by two hypotheses concerning ToP.
4.1 Datasets
We begin by collating a collection of graph datasets. Statistics are available in the Supplemental Material in Table 3. Here we provide qualitative detail as well as any downstream task information:
@@ 4.1 Datasets
The full neural connectome of a fruit fly larvae. As the multi-graph is dense at originally 548,000 edges, we include a single edge between neurons if there are more than two synapses between them. Used only in pre-training.
@@
A single graph of page-page connections on Facebook. Used both in pre-training and downstream transfer. On a graph level the downstream task is predicting average clustering, and we also evaluate transfer to node classification.
@@
Ego networks from the streaming platform Twitch, used both in pre-training and downstream transfer, with a downstream binary classification task.
@@
The road network of Pennsylvania. Used both in pre-training and downstream transfer, with the task set to be predicting graph diameter normalised by graph size.
@@
molesol, molfreesolv, mollipo, molclintox, molbbbp, molbace, molhiv. Smaller datasets sampled from MoleculeNet, with single-target classification or regression downstream tasks. These are used only in downstream transfer.
@@
Trees, of a randomly sampled maximum depth, and a randomly fixed probability of branching at each level, with a downstream task of predicting depth. Used only in downstream evaluation.
@@
Erdos-Renyi (ER) graphs [7], with a number of nodes and edge probability randomly sampled, and predicting edge probabilities as a downstream task. Used only in downstream evaluation.
@@
Small ER subgraphs, with a set intra-subgraph and random inter-subgraph connection probability, with inter-subgraph connectivity as target. Used only in downstream evaluation.
As the Facebook, Cora, Roads and Fly Brain datasets consist of a single graph, we employ exploration sampling to construct datasets of many smaller graphs. We use a randomly selected exploration sampler [^1] Using the littleballoffur Python library [27]. for each sample, which in turn samples a random number of nodes in the range $24\leq|V|\leq 96$ from the source graph. The use of a selection of exploration samplers should further ensure that a large variety of graphs and topological features are included in the datasets.
Figure 2: A UMAP embedding of encodings from each model, as well as an untrained model. The untrained and Chem models are noticeably more fragmented than the Social and All models. In turn the Social model is more fragmented than the All model. In the All model embedding, molecules (primarily dark blue) are all in the same region, furthest from the Twitch Ego networks (lilac), with the Cora (brown) and Facebook Page-Page data (pink) “filling the gap”.
We take the downstream task on each validation set where available and add their relevant scores (MSE and $1-$ AUROC) with a linear model as a crude monitor for the models’ representational abilities. We conduct a limited Bayesian hyperparameter sweep. The heuristic used is addition of evaluation scores on each dataset weighted by the number of samples in that dataset.
4.2 Models
We perform a limited ablation study by evaluating pre-training five models on different subsets of the amalagamated dataset: 1) GIN, an untrained GIN model of identical architecture. 2) ToP-Chem, an AD-GCL model fit only on the molecular molpcba dataset. 3) ToP-Social, an AD-GCL model fit only on non-molecular datasets (Cora, Facebook, Fly Brain, Roads, Twitch). 4) ToP-All, an AD-GCL model fit on all the training data. 5) All (CL), a model trained with the loss function from You et al. [38] using random edge dropping ($\lambda=0.2$).
Models during pre-training, unless specified, are six GIN [37] layers with single linear layers as input and output. The GIN model acts as our supervised baseline (Hypothesis 1). The ToP-Chem/ToP-Social models indicate the impact of training data diversity on downstream transfer performance, particularly on our downstream molecular datasets (Hypothesis 2). The ToP-All (CL) model shows the benefits of adversarial augmentations.
Following the results of the hyperparameter sweep, we train each model for 100 epochs with a batch size of 512, learning rates of 0.001, embedding and projection dimensions of 300 and a regularisation constant of 0.2. Figure 2 shows a UMAP embedding of encodings of the validation datasets for an untrained model and the pre-trained models. A larger version can be found in the Supplemental Material.
Pre-training was conducted on a workstation with an NVIDIA A6000, using around 16GB of video memory capacity. Pre-training the ToP-All model took a little more than eight hours. Transfer results were produced on consumer laptops, primarily with an Apple M1 processor, and on GPU use around 3GB of video memory with batch sizes of 128.
@@ 4.3 Transfer
Here we apply our pre-trained models to various downstream tasks through transfer learning. This follows stages of abstraction away from ToP pre-training: first we fine tune on datasets without features, then on datasets with features included, and finally on node and edge-level tasks.
Table 1: Scores for fine-tuning on each validation dataset. For the classification datasets we report AUROC and for regression datasets RMSE. Bold text indicates a result beats the unsupervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. Checkmarks ✓indicate that the ToP-All model significantly out-performs the GIN baseline at $p<0.01$. ‘Best’ indicates the proportion of datasets over which a model has superior performance, and ‘Positive’ the proportion of datasets over which performance is superior to the fully-supervised baseline.
GIN (Baseline) All (CL) ToP-Chem ToP-Social ToP-All
@@
Features Excluded We replace the output head with a simple two-layer MLP, with a single output node. We then perform 10 fine-tuning runs on each validation dataset, with and without features, for each pre-trained model. Table 1 presents results of transfer and fine-tuning without features.
@@
The ToP-All model consistently out-performs the other models under fine-tuning on the majority of datasets. At $p\leq 0.01$ the ToP-All model significantly out-performs the GIN baseline on 77% of datasets. When performance is not significantly better, it is never significantly worse. Notably the ToP-All model always out-performs the All (CL) model, indicating the merits of using a learnt augmentation scheme in this multi-domain setting. The standard deviation of the results on each dataset are lower than those for the non-pretrained model and the domain-specific pre-trained models. In Figure LABEL:fig:training-examples in the supplemental material we show validation loss on the several datasets during fine-tuning for the ToP-All model and an un-pre-trained GIN with the same architecture. Here the advantages of pre-training are clear: the pre-trained model has consistently lower validation loss, seems to plateau later and at a lower loss, and has a lower deviation in that loss than the baseline GIN.
Pre-training only on molecules (‘Chem’ data subsets) results at-best on-par transfer compared to the supervised GIN baseline, with negative transfer on the majority of datasets. Pre-training only on non-molecules, in comparison, results in positive transfer on the large majority of datasets and pre-training methods. At $p\leq 0.01$ the ToP-Social model significantly out-performs the ToP-Chem model on the majority of datasets. As before, where performance is not significantly better, it is never significantly worse. We employ an ablation study over the non-molecular pre-training datasets (see Supplemental Material) to ensure that this is not a single artifact result. This study corroborates the results presented in Table 1, with the Chem model consistently out-performed by non-molecular models.
@@
We find that under pre-training with random edge dropping performance actually drops when including molecular data (‘Social’ vs ‘All’). This suggests that AD-GCL is able to differentiate between molecules and non-molecules and applies different augmentations to these domains. Nonetheless, we show that using a variety of non-molecular graphs in-general provides a more solid pre-training foundation than including molecules for both methods of augmentation.
@@
Features Included By swapping the input head of the pre-trained model, as seen in Figure 1, arbitrary node and edge features can be included during transfer. To this end we replace the input head of our model with a three layer Multi-Layer Perceptron (MLP) for both node and edge features, then perform fine-tuning runs on our molecular datasets. Table 2 presents the results of this transfer and fine-tuning. These tasks are assumed to require node and edge features to obtain good performance, and here positive transfer rates are similar to when no features are included, despite ToP’s exclusion of features during pre-training. The ToP-All and ToP-Social models again demonstrate consistent positive transfer, and on most datasets their performance improves compared to when features were excluded. The same results are presented for different GNN backbones in Table 9 in the Supplemental Material, but perform no optimisation for these models, using hyperparameters from the GIN models here for the GAT and GCN models presented. As such these results are included only for completeness.

### v4

@@ 3.1 Hypothesis 1: Excluding Features
Pre-training on multiple domains with node and edge features excluded can lead to consistent and positive transfer compared to fully-supervised, task-specific models. The motivating concept behind ToP is that as graph learning requires learning on both structural and feature levels, pre-training on only structure can lead to useful knowledge transfer. As such, a model that hasn’t undergone such pre-training should perform worse on downstream tasks when applied under the same supervised training conditions.
Supporting Hypothesis 1 means producing a model that has learnt useful representations of graphs in general. Our first step to achieving this is to construct a dataset that covers a variety of graph domains. Such a dataset should contain graphs that present a large diversity of multi-node patterns. Our approach is simple: include graphs from several domains and assume that there is adequate coverage of the multi-node patterns that occur in real-world graphs for a given set of downstream tasks. The same approach has been taken in the domains of text [3] and images [34], where enormous numbers of data samples are scraped from the web for generative pre-training. This relies on two assumptions 1) this dataset adequately covers the relevant space for downstream transfer, and 2) a single model can usefully represent such a space. These assumptions, as in other fields, are justified through performance metrics on downstream tasks on seen and unseen domains.
@@ 3.2 Hypothesis 2: Sample Diversity
With the exclusion of node and edge features, out-of-domain pre-training can have greater performance benefits than in-domain pre-training across downstream tasks. If structure-only learning is useful, as proposed by Hypothesis 1, it follows that out-of-domain of pre-training data with greater diversity may be more useful than in-domain. This relies on the assumption that a given encoder is learning to represent local node patterns, and as diverse data presents a wider range of such patterns, the model must learn more expressively how to represent such patterns.
From Hypothesis 2 we expect that for a given task models trained on a more diverse pre-training domain can bring better transfer than a model trained on the target domain. We do not argue that this is always true, only that in some settings out-of-domain samples can be more useful for pre-training than in-domain. This has clear parallels to the performance-by-scale approach taken in other fields [38, 9, 30]. To this end we compare two models, the first trained on a dataset from a single domain, as is the current practice. This single domain matches that of the downstream target dataset(s). The second model is instead trained on the rest of the data while excluding the target domain.
@@
3.3 Models for Evaluation
In order to address these hypotheses, we apply ToP over five principal different combinations of pre-training data and augmentation scheme.
GIN
An untrained GIN model of identical architecture to the other methods. This serves to benchmark what constitutes positive transfer.
ToP-Chem
An AD-GCL model fit only on molecules. This constitutes an in-domain model, albeit in the absence of domain features.
ToP-Social
An AD-GCL model fit only on non-molecules. This serves as a contrast to ToP-Chem, allowing us to assess the benefits of varied out-of-domain pre-training in comparison to in-domain pre-training.
ToP-All
An AD-GCL model fit on all the training data. This is our primary measure of the effectiveness of ToP.
ToP-All (CL)
A model trained across the whole amalgamated dataset with the loss function from You et al [54] using random edge dropping ($\lambda=0.2$). This allows us to assess the impact of adversarial augmentations in ToP pre-training.
Models during pre-training, unless specified, are six GIN [53] layers with single linear layers as input and output. The GIN model acts as our supervised baseline (Hypothesis 1 in Section 4). The ToP-Chem/ToP-Social models indicate the impact of training data diversity on downstream transfer performance, particularly on our downstream molecular datasets (Hypothesis 2). The ToP-All (CL) model shows the impact of adversarial augmentations.
Figure 3: UMAP embeddings of encodings from each model, as well as an untrained GIN model. Triangular markers show molecular graphs, and circles non-molecules. We plot the centroid of each dataset for added clarity. Qualitatively the untrained and ToP-Chem models are noticeably more fragmented than the ToP-Social and ToP-All models. In turn the ToP-Social model is more fragmented than the ToP-All model.
@@ 4 Experiments
4.1 Datasets for Evaluation
Evaluating ToP requires collecting datasets of graphs from multiple domains. A portion of these datasets should have downstream tasks, the useful information for which relies on both structure and features, with varied weighting to one or the other. To this end we curate the following datasets, with Table 6 in the Appendix providing dataset statistics:
Pre-Training Only
@@ 4.1 Datasets for Evaluation
The full neural connectome of a fruit fly larvae. As the original multi-graph is dense at 548,000 edges, we include a single edge between neurons if there are more than two synapses between them. Used only in pre-training.
Pre-Training and Evaluation
@@
A single graph of page-page connections on Facebook. Used both in pre-training and downstream transfer. On a graph level the downstream task is predicting average clustering, and we also evaluate transfer to node classification. This task does not use feature information.
@@
Ego networks from the streaming platform Twitch, used both in pre-training and downstream transfer, with a downstream binary classification task. This task does not use feature information.
@@
The road network of Pennsylvania. Used both in pre-training and downstream transfer, with the task set to be predicting graph diameter normalised by graph size. This task does not use feature information.
Evaluation Only
@@
molesol, molfreesolv, mollipo, molclintox, molbbbp, molbace, molhiv. Datasets sampled from MoleculeNet, with single-target classification or regression downstream tasks. These are used only in downstream transfer. These benchmarks have pre-prepared node and edge features, allowing us to measure the extent to which ToP models can integrate feature information alongside structural. We source train/validation/test splits from the Open Graph Benchmark [19].
@@
Trees, of a randomly sampled maximum depth, and a randomly fixed probability of branching at each level, with a downstream task of predicting depth. Used only in downstream evaluation. This task does not use feature information.
@@
Erdos-Renyi (ER) graphs [11], with a number of nodes and edge probability randomly sampled, and predicting edge probabilities as a downstream task. Used only in downstream evaluation. This task does not use feature information.
@@
Small ER subgraphs, with a set intra-subgraph and random inter-subgraph connection probability, with inter-subgraph connectivity as target. Used only in downstream evaluation. This task does not use feature information.
Datasets from Single Graphs
As the Facebook, Cora, Roads and Fly Brain datasets consist of a single graph, we employ exploration sampling to construct datasets of many smaller graphs. We use a randomly selected exploration sampler for each sample [36], which in turn samples a random number of nodes in the range $24\leq|V|\leq 96$ from the source graph. The use of a selection of exploration samplers should further ensure that a large variety of graphs and structural patterns are included in the datasets.
Hyperparameters
We take the downstream task on each validation set where available and add their relevant scores (MSE and $1-$ AUROC) with a linear model as a crude monitor for the models’ representational abilities during pre-training. 5,000 samples are used for non-molecular datasets. We conduct a limited Bayesian hyperparameter sweep. The heuristic used is addition of evaluation scores on each dataset weighted by the number of samples in that dataset. Following the results of the hyperparameter sweep, we train each model for 100 epochs with a batch size of 512, learning rates of 0,001, embedding and projection dimensions of 300 and a regularisation constant of 0,2. This is very close to the default AD-GCL parameters. Figure 3 shows UMAP embeddings of encodings of the validation datasets for our models. Larger versions for the four ToP models can be found in the Appendix.
Compute
Pre-training was conducted on a workstation with an NVIDIA A6000, using around 16GB of VRAM capacity. Pre-training the ToP-All model took a little more than eight hours. Transfer results were produced on consumer laptops, primarily with an Apple M1 processor, and on GPU use around 3GB of video memory with batch sizes of 128.
@@ 4.2 Transfer
Here we apply our pre-trained models to various downstream tasks through transfer learning. This follows stages of abstraction away from ToP pre-training: first we fine tune on datasets without features, then on datasets with features included, and then compare ToP models to reasonable benchmarks. Finally we report performance on node and edge-level tasks.
Table 1: Scores for fine-tuning on each validation dataset. Molecular benchmarks here do not include node or edge features. For the classification datasets we report AUROC and for regression datasets RMSE. Bold text indicates a result beats the supervised (GIN) baseline. Underlined text indicates a superior result, down to three significant figures. Checkmarks ✓indicate that the ToP-All model significantly out-performs the GIN baseline at $p<0.01$. ‘Best’ indicates the proportion of datasets over which a model has superior performance, and ‘Positive’ the proportion of datasets over which performance is superior to the fully-supervised baseline.
GIN ToP-All (CL) ToP-Chem ToP-Social ToP-All
@@ 4.2.1 Features Excluded
We replace the output head with a simple two-layer MLP. We then perform 10 fine-tuning runs of 50 epochs on each validation dataset, with and without features, for each pre-trained model. Table 1 presents results of transfer and fine-tuning without features.
@@
The ToP-All model consistently out-performs the other models under fine-tuning on the majority of datasets. At $p\leq 0.01$ the ToP-All model significantly out-performs the GIN baseline on 77% of datasets. When performance is not significantly better, it is never significantly worse. The standard deviation of the results on each dataset are lower than those for the non-pretrained model and the domain-specific pre-trained models. In Figure LABEL:fig:training-examples in the Appendix we show validation loss on the several datasets during fine-tuning for the ToP-All model and an un-pre-trained GIN with the same architecture. Here the advantages of pre-training are clear: the pre-trained model has consistently lower validation loss, seems to plateau later and at a lower loss, and has a lower deviation in that loss than the baseline GIN.
On our evaluation-only datasets (Trees, Community, Random), on which no models are pre-trained, we see markedly better performance from models that are pre-trained on a range of domains. In particular the ToP-All and ToP-Social models drastically out-perform our supervised baseline and ToP-Chem. On the Trees dataset, which intuitively and on aggregate measures is the closest of the evaluation-only datasets to molecular data, ToP-Chem also has negative transfer.
@@
Pre-training only on molecules (‘Chem’ data subsets) results at-best on-par transfer compared to the supervised GIN baseline, with negative transfer on the majority of datasets. Pre-training only on non-molecules, in comparison, results in positive transfer on the large majority of datasets and pre-training methods. At $p\leq 0.01$ the ToP-Social model significantly out-performs the ToP-Chem model on the majority of datasets. As before, where performance is not significantly better, it is never significantly worse. This study corroborates the results presented in Table 1, with the Chem model consistently out-performed by non-molecular models. We find that under pre-training with random edge dropping performance actually drops when including molecular data (‘Social’ vs ‘All’). This suggests that AD-GCL is able to differentiate between molecules and non-molecules and applies different augmentations to these domains. Nonetheless, we show that using a variety of non-molecular graphs in-general provides a more solid pre-training foundation than including molecules for both methods of augmentation.
4.2.2 Features Included
We replace the input head of our model with a three layer Multi-Layer Perceptron (MLP) for both node and edge features, then perform fine-tuning runs on our molecular datasets. Table 2 presents the results of this transfer and fine-tuning. These tasks are assumed to require node and edge features to obtain good performance, and here positive transfer rates are similar to when no features are included. This shows that the focus on structure during ToP pre-training does not inhibit the use of feature-based information in downstream transfer. The ToP-All and ToP-Social models again demonstrate consistent positive transfer, and on most datasets their performance improves compared to when features were excluded. The same results are presented for different GNN backbones in Table 11 in the Appendix, but we perform no optimisation for these models, using hyperparameters from the GIN models here for the GAT and GCN models presented. As such these results are included only for completeness.
@@ 4.2.2 Features Included
4.3 Comparison to Other Methods
The selection of valid benchmarks from other works is limited. We start by splitting existing works into two broad categories, generalists that function on any graph domain, and specialists that are designed for a specific domain of graph data. The selection of specialist graph models in chemistry is broad, but not necessarily a fair comparison, as these works do not share the same fundamental aims as ToP. They are something of a best-case for pre-training: abundant pre-training data, without missing features. In the following sections we include specialist GNNs, all pre-trained with constrastive learning, as well as specialist LLMs designed for molecular tasks. We don’t expect ToP models to out-perform these specialists. Instead, we include these results to evaluate how our multi-domain pre-training framework compares to the current state of the art within a domain.
The selection of generalist models is far more limited. GCC is, to the best of the author’s knowledge, the only other generalist GNN work that does not rely on some level of feature transformation using LLMs. This carries the significant caveat that the original work completely eliminates features, and the information they carry, from downstream transfer. We extend the original GCC work somewhat, swapping the input and output heads as with our own models, in order to provide fair comparison against ToP.
The use of LLMs to unify the feature input space is we argue, a level of specialism in-and-of itself, as these datasets cannot be immediately applied to new domains, instead requiring hand-design of a textual schema. LLMs themselves have been argued to be generalists [28], in that any graph structure can function as input given some encoding schema.
4.3.1 Molecular Tasks
In Table 3 we compare ToP to other pre-training schemes across molecular tasks. Results are sourced from the original works. These methods are divided into specialist (domain-specific) and generalist (domain-agnostic) approaches. For domain-specific approaches we include InfoGraph [51], GraphCL [17] and AD-GCL [43], contrastive learning approaches. These models are pre-trained with all features included, and hyper-parameters optimised, over Zinc, a dataset of around 2M molecules. We also include MolXPT [27], InstructMol [4] and ChemDFM [58], LLM-based approaches.
Table 3: Our results for the ToP-All and ToP-Social models presented alongside the same results from other pre-training works. Alongside dataset names we include the total size of the dataset. Results are ROC-AUC, and other than for ToP and GCC, taken verbatim from the respective work. Where standard deviations or errors are not present, they were also not present in the respective work. The table is organised by specialism and type (Specialist GNN, Specialist with LLM, Generalist LLM, Generalist GNN). ($n$B) indicates $x$ billion parameters. Non-LLMs are GINs with 6 layers and in the region of 7 million parameters. Bold text indicates a superior result within a type of model.
clintox (1329) bbbp (1835) bace (1361) hiv (37014)
InfoGraph (2.4M) 0.70 $\pm$ 0.03 0.69 $\pm$ 0.01 0.76 $\pm$ 0.02 0.76 $\pm$ 0.02
[51]
GraphCL (2.4M) 0.76 $\pm$ 0.03 0.70 $\pm$ 0.01 0.75 $\pm$ 0.01 0.78 $\pm$ 0.01
[17]
AD-GCL (2.4M) 0.80 $\pm$ 0.04 0.70 $\pm$ 0.01 0.79 $\pm$ 0.01 0.78 $\pm$ 0.01
[43]
MolXPT (350M) 0.95 $\pm$ 0.00 0.80 $\pm$ 0.01 0.88 $\pm$ 0.01 0.78 $\pm$ 0.00
[27]
InstructMol (7B) 0.70 0.82 0.69
[4]
ChemDFM (13B) 0.90 0.67 0.78 0.74
[58]
GPT-4 ($>$200B [^1] Best-guess speculation, as OpenAI is closed-source) 0.52 0.62 0.63 0.66
[31]
LLaMa-2 (13B) 0.46 0.60 0.26 0.29
[45]
Galactica (30B) $\dagger$ 0.82 0.60 0.73 0.76
[44]
GCC (untrained) (200k) 0.50 $\pm$ 0.03 0.59 $\pm$ 0.01 0.44 $\pm$ 0.04 0.46 $\pm$ 0.01
GCC (200k) 0.40 $\pm$ 0.07 0.52 $\pm$ 0.02 0.50 $\pm$ 0.04 0.52 $\pm$ 0.01
GCC+Feat (untrained) 0.79 $\pm$ 0.04 0.64 $\pm$ 0.02 0.70 $\pm$ 0.05 0.70 $\pm$ 0.03
GCC+Feat 0.50 $\pm$ 0.04 0.61 $\pm$ 0.03 0.70 $\pm$ 0.1 0.68 $\pm$ 0.05
ToP-Social (2.4M) 0.86 $\pm$ 0.02 0.62 $\pm$ 0.03 0.70 $\pm$ 0.05 0.62 $\pm$ 0.11
ToP-All 0.84 $\pm$ 0.05 0.67 $\pm$ 0.02 0.72 $\pm$ 0.05 0.57 $\pm$ 0.13
We also include results from other domain-agnostic graph models. For LLM-based approaches these chemical tasks these results are sourced the ChemDFM work [58]. Performing molecular tasks with these models relies on the same assumptions as the chemistry-specific models, but without fine-tuning specifically on chemical literature [^2] Galactica is trained on scientific literature, including but not limited to chemistry.
Finally we include GCC, to our knowledge the only other framework or approach for producing generalised graph models without relying on LLMs, albeit at the cost of feature inclusion for downstream tasks. We fine-tune the two pre-trained encoders packaged with GCC [^3] Our GCC version, updated to current packages, is available at https://github.com/alexodavies/GCC-Updated. Given that GCC does not provide a method for re-introducing domain features, we take the same approach as with our ToP models, swapping the projection heads for nodes and edges (GCC+Feat). We also include GCC with positional encodings as features, as in the original work (GCC). Alongside the pre-trained GCC models we also include results for an equivalent but un-pre-trained model. Where applicable we use the same hyper-parameters as for our fine-tuning of ToP models. In practise this is a batch size of 512 and a learning rate of 0,001.
We find that AD-GCL performs most strongly among the specialist GNN models, MolXPT strongest among the specialist LLMs, Galactica among the generalist LLMs, and ToP-All among the generalist GNNs. Among specialists MolXPT performs most strongly. Among generalists Galactica and our ToP models perform comparably, though both strongly out-perform other models in both classes.
GCC models show consistent negative transfer, both when using their original positional encoding features and when we apply the ToP process for swapping input heads. On the much larger molhiv dataset the smaller (200k parameters vs 7M) GCC models reach greater performance within our 50 epoch fine-tuning runs. Performance from GCC models is consistently better when we apply techniques from this work to include specialist features in place of positional encodings using the ToP approach. As the GCC encoders are much smaller than out ToP encoders, we suspect that smaller models trained with ToP would reach comparable performance.
4.3.2 Non-Molecular Tasks
We apply two recent LLMs, ChatGPT-4o-mini [31] and Llama-3.2-Instruct (3B) [10], across our validation-only datasets (random, community, trees), as well as the Twitch Egos dataset. We use these smaller versions from their respecitve LLM families to reduce computational cost, which is still drastically higher than our GNN-only approach. From here we denote Llama-3.2-Instruct (3B) as Llama-3. These tasks are chosen as they do not rely on node or edge features, meaning that learning and performance directly follows the use of structural information. We also include GCC with positional encodings as features, as in the original work.
LLM Prompting
Using LLMs for arbitrary graph structures requires defining a schema for natural language representations of graphs. We opt for the ‘incident’ encoding style from Fatemi et al [12]. In this format a graph is encoded in the style nodes: $(1,2,3,...)$ edges: $(1\leftrightarrow 2,2\leftrightarrow 3,...)$. We define prompts for each dataset, aimed to both provide contextual information and make tasks clear, which can be found in Listing 2 in the Appendix. We include further prompting text in an attempt to ensure uniformity across responses: ‘Return a number, with no other text or filler. Do not explain your working. Answer in the format THE ANSWER IS X.’ The prompt structure is then {task description + extra prompt information + graph information}. We show an example in Listing 1, and an example response from Llama-3 in the Appendix in Listing 3.
Listing 1: An example prompt for LLMs on the Trees dataset.
question: This is a Random graph. What is the connection probability between two nodes?
Return a number, with no other text or filler. Do not explain your working. Answer in the format THE ANSWER IS X.
graph: nodes: 0, 1, 2, 3, ...
edges: 0 <-> 23, 0 <-> 17, 1 <-> 5, 1 <-> 16, ...
Table 4: Our results for the ToP-All and ToP-Social models presented alongside LLMs and GCC applied to structure-only datasets. Random, Community and Trees are unseen data for all models. Twitch Egos is AUC-ROC, other datasets are RMSE. Due to cost constraints we do not produce error margins for LLM models.
Twitch Egos Random Community Trees
GPT-4o-mini (8B [^4] Best-guess speculation, as OpenAI is closed-source) 0.50 0.30 0.37 0.28
LLaMa-3.2 (3B) 0.46 0.12 0.10 0.42
GCC (untrained) 0.66 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.01 0.12 $\pm$ 0.01
GCC (200k) 0.66 $\pm$ 0.01 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00
ToP-Social (7M) 0.74 $\pm$ 0.01 0.04 $\pm$ 0.00 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00
ToP-All 0.79 $\pm$ 0.01 0.04 $\pm$ 0.00 0.02 $\pm$ 0.00 0.12 $\pm$ 0.00
As can be seen in Listing 3, models frequently ignored instructions to return only a number, and instead generated large amounts of text. To extract numbers from generated text, we design a simple process, shown in Appendix in Listing 4.
Findings
LLMs, across all tasks, perform worse than ToP or GCC models. This is despite three of our datasets consisting of basic counting tasks, meaning that they are possible to perform in natural language. Larger LLMs might perform better, but given that these models are already many orders of magnitude larger than ToP or GCC models, we view the LLMs used here as fair benchmarks. On the three synthetic datasets (Random, Community, Trees), with their simple structure-based tasks, ToP models and GCC reach essentially the same performance.
That said, a non-pretrained version of the small GCC encoder produces almost exactly the same performance as the pretrained encoder. This indicates that in this setting GCC pre-training has minimal positive transfer benefits. Also worth noting is that the positional encodings GCC uses as node and edge features are explicitly useful for these tasks. ToP models reach the same performance with no such advantage.
On the real-world Twitch Egos dataset ToP models strongly out-perform the LLMs and GCC. A caveatte is that ToP-Social and ToP-All both have graphs from this dataset in their pre-training data. In Section 4.4.1 we present results for ToP models with the Twitch Ego networks excluded from pre-training data, which under fine-tuning had an ROC-AUC performance of 0.73 $\pm$ 0.03, still significantly better than GCC.
4.3.3 Further Transfer Results